# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v2

@@ 3.3.1 Enhancement
In low-light conditions, scenes often exhibit low contrast, leading to focusing difficulties or the need for long exposures, which can result in blurred images and videos. To address this, LEDNet [249] has introduced a synthetic dataset for such scenarios and incorporated a learnable non-linear activation function within the network to enhance feature intensities. Meanwhile, SNR-Aware [94] estimates spatial-varying Signal-to-Noise Ratio (SNR) maps and proposes local and global learning branches using ResNet and transformer architectures, respectively. NeRCo [104] address low-light problem with INR, which unifies the diverse degradation factors of real-world scenes with a controllable fitting function. Diffusion models (DMs) have also become popular choices for low-light image enhancement [100, 101, 102]. Diff-Retinex [101] formulates the low-light image enhancement problem into Retinex decomposition, and employs multi-path generative diffusion networks to reconstruct the normal-light Retinex probability distribution. A recent state-of-the-art approach presented in [102] decomposes images into high and low frequencies using wavelet transform. High frequencies are enhanced using a transformer-based pipeline, while the low frequencies undergo a diffusion process. This method achieves nearly 2.8dB improvement over the state-of-the-art transformer-based approach, e.g. LLFormer [96], and significantly better than INR-based method, NeRCo [104], on a real low-light image benchmarking dataset. The technique has been extended for video enhancement in [103]. The output of the enhancement typically depends on user preferences. This has been viewed as a one-to-many inverse problem, with attempts to solve it using Bayesian approaches. For example, a Bayesian Enhancement Model (BEM) [250] incorporates Bayesian Neural Networks (BNNs) to capture data uncertainty and produce diverse outputs. The method can be used with Transformers or Mamba as the architecture backbone. CLE Diffusion [123] offers lighting that is editable with user-friendly region controllability.
@@
Regarding video enhancement, transformer and DMs are still in their early stages. STA-SUNet [97] has demonstrated that using transformers for low-light video enhancement outperforms CNN-based methods [251]. It is important to note that low-light enhancement is subjective. While most training datasets use normal lighting conditions as ground truth [252], the enhanced images and videos may alter the mood and tone of the content. Therefore, the tools for creative industries should be adjustable, not only for entire images and videos but also adaptive to specific areas and content.
@@ 3.3.2 Style transfer
StyTr2 [105] is the first transformer-based method for style transfer, applying content as a query and style as a key of attention. InST [108] utilizes Stable Diffusion Models as the generative backbone and introduces an attention-based textual inversion module to learn the description of the content. StableVideo [109] uses a text prompt to describe the desired appearance of the output, transforming the input video to have a new look based on a diffusion model. For instance, a video of a white car driving in summer can be altered to show a red car driving in winter. A large pre-trained DM is employed in [107], where the style is injected to manipulate the self-attention of the decoder. To deal with the disharmonious color, they propose an adaptive instance normalization. Implicit Neural Representation (INR) is combined with vision transformers in [106] for generalizable style transfer; however, the results are not very impressive. The method proposed in [110] employs MLPs to map image coordinates to colors of the new style image, controlled by features from the content input and style input.
@@ 3.3.3 Upscaling imagery: super-resolution (SR)
The first use of a transformer, called ESRT, was for capturing long-term dependencies, such as repeating patterns in buildings. This was done in the feature domain extracted by a lightweight CNN module [112], outperforming those that use only CNNs. Since then, most SISR methods have been based on transformers. The Hybrid Attention Transformer (HAT) [114] was introduced, which improves the SR quality over ESRT by more than 2dB when upscaling 2$\times$-4$\times$. However, the NTIRE 2023 Real-Time Super-Resolution Challenge [253] showed that the winner, Bicubic++ [254], uses only convolutional layers and achieves the fastest speed at 1.17ms in upscaling 720p to 4K images.
@@
For video SR, numerous methods have emerged as part of a unified enhancement framework, as discussed in the previous section. One of the pioneering works to incorporate transformers specifically for video SR tasks is the Trajectory-aware Transformer for Video Super-Resolution (TTVSR) [113]. Although the results are slightly inferior to those of BasicVSR++ [256], which employs CNN and was introduced around the same time, both methods significantly enhance detail and sharpness compared to previous approaches, albeit not in real time. To address this limitation, the Deformable Attention Pyramid [257] has been introduced, offering slightly lower quality but a speed-up of over 3$\times$. Recently, Adobe announced their VideoGigaGAN [117], which can perform 8$\times$ upsampling. This is achieved by adding flow estimation and temporal self-attention to the GigaGAN upsampler [116], which is primarily used for image SR, and text-to-image synthesis.
@@ 3.3.4 Restoration
The pioneering transformer-based method for image restoration, SwinIR [111], employs several concatenated Swin Transformer blocks [16]. SwinIR surpasses state-of-the-art CNN-based methods proposed up to the year 2021 in super-resolution and denoising tasks. The model is smaller and reconstructs fine details more effectively. Other two popular approaches that emerged in the same timeframe are Uformer [124] and Restormer [125]. Both incorporate Transformer blocks into hierarchical encoder-decoder networks, employing skip connections similar to those in U-Net. Their objective was to restore noisy images, sharpen blurry images, and remove rain. The networks focused on predicting the residual $R$ and obtaining the restored image $\hat{x}$ through $\hat{x}=y+R$. While their performance is very similar, Restormer has half the parameters of Uformer. More recent, GRL by Li et al. [115] exploits a hierarchy of features in a global, regional, and local range using different ways to compute self-attentions as an image often show similarity within itself in different scales and areas. GRL outperforms SwinIR and Restormer. Additionally, Fei et al. introduced the Generative Diffusion Prior [122] for unsupervised learning, aiming to model posterior distributions for image restoration and enhancement.
@@
For audio restoration, most software discussed in Section 3.1.2 offers tools for enhancing audio quality, such as eliminating background noise, echo, microphone rumble, and occasionally room reverberation, which have been well-established even before the advent of deep learning. There have been efforts to utilize AI for learning global contextual information to aid in the removal of unwanted sounds, leading to better final quality [128]. The latest advancements in this domain are primarily focused on addressing issues where significant portions of the audio data are missing. For instance, Moliner et al. [119] tackle problems such as audio bandwidth extension, inpainting, and declipping by treating them as inverse problems using a diffusion model.
@@
i) Deblurring: A lightweight deep CNN model was recently proposed in [258], where a new discriminative temporal feature fusion has been introduced to select the most useful spatial and temporal features from adjacent frames. Feature propagation along the video is done in the wavelet domain. The deblurring performance is comparable to RVRT [95], but it is 5 times faster. DaBiT [127] mitigates focal blur content with depth information and applies SR for further enhancing fine details. Note that not only in software, but AI technologies have also been integrated into hardware. This includes autofocus, which is crucial for capturing sharp images of subjects, especially in dynamic environments where manual adjustments are impractical due to rapid movement. AI-driven autofocus methods have emerged, often tailored for specific camera hardware. For instance, Choi et al. proposed an autofocus model optimized for dual-pixel Canon cameras [259]. Additionally, Yang et al. investigated the correlation between language input and blur map estimation, utilizing semantic cues to enhance autofocus performance [126]. Remarkably, their model achieves comparable results to previous state-of-the-art methods while being more lightweight [260]. Autofocus could be used in conjunction with real-time object tracking (see Section 3.4.3) to produce desirable sharpness for moving objects in the video.
@@
ii) Denoising: SUNet [18] applies Swin transformer blocks combined in a UNet-like architecture. Denoising with diffusion models (DMs) [134] has been proposed by diffusing with estimated noise that is closer to real-world noise rather than Gaussian noise, achieving better performance than SwinIR [111] and Uformer [124]. INR with complex Gabor wavelets as activation functions show promising denoising results [44].
@@
iii) Dehazing: Vision transformers for single image dehazing was proposed in DehazeFormer [129]. Similar to SUNet, it is UNet-like architecture, but introduced Rescale Layer Normalization for better suit on improving contrast. For video dehazing, Xu et al. [130] introduced a recurrent multi-range scene radiance recovery module with the space-time deformable attention. They also employs physics prior to inform haze attenuation. This method outperforms DehazeFormer by approximately 1dB.
@@
iv) Mitigating atmospheric turbulence: Similar to dehazing, physics-inspired models have been widely developed to remove turbulence distortion [136, 137], while complex-valued CNNs have been proposed to exploit phase information [261]. There was also an attempt to use instance normalization (INR) to solve this problem, offering tile and blur correction [137]. However, diffusion models outperform on a single image [135], and transformer-based methods remain state-of-the-art for restoring videos [132, 133]. A recent review can be found in [262].
@@ 3.3.5 Inpainting
A survey of learning-based image and video inpainting, including CNN, VAE, GAN, transformers and diffusion models can be found in [265].
@@ 3.3.6 Image Fusion
Image fusion involves merging multiple images from either the same source (such as varying focal points or exposures) or different modalities (like visible and infrared cameras) into a single image. This process integrates complementary information from the various images to enhance overall quality, improve interpretation, and increase the usability of the final image.
Transformers and CNNs have been combined to extract global and local information, respectively. Most methods use CNNs for feature extraction, with transformers operating in the latent space [143, 144]. Notable methods include SwinFusion [143], which utilizes a self-attention-based intra-domain fusion unit and a cross-attention-based inter-domain fusion unit to achieve multi-modal and digital photography image fusion. Transformer-based image fusion has also been applied to downstream tasks like segmentation [145], achieving superior results by leveraging the additional information.
DDFM, the first diffusion model-based image fusion method, has been introduced, using loss functions to estimate noise in the reverse process by combining multiple inputs [146]. The expectation-maximization (EM) algorithm is integrated to estimate the noise distribution parameters, resulting in sharper images compared to traditional DDPM. For an in-depth review, refer to recent works in [266, 267].
@@ 3.3.7 Editing and Visual Special Effects (VFX)
VFX aims to create and/or manipulate imagery outside the context of a live-action shot in filmmaking and video production. When adding objects, scenes, and effects into traditional photographic videos, generative AI has obviously become a great tool, but some manual operations are still required. For example, in After Effects (EA) [^40] https://www.adobe.com/uk/products/aftereffects.html, the user selects the area where the object will be added and uses text prompts to describe such object. Subsequently, with the current EA version, the user will need to apply motion tracking so the generated object is moved accordingly.
AI technologies can upscale, enhance, and restore low-quality or old footage. For example, standard definition videos can be converted to high definition or even 4K quality without traditional manual remastering processes. This is particularly useful for remastering old movies or enhancing visual details in scenes. Generative AI has also simplified and accelerated automated processes, such as rotoscoping [268], an animation technique where animators trace over motion picture footage frame by frame to create realistic action. AI models can accurately detect and segment objects and characters in video frames, significantly speeding up the post-production process. Additionally, AI assists in the rapid creation of 3D models from 2D images and generates realistic animations from minimal input data, facilitating complex human motions and synchronized facial expressions to voiceovers. One restriction is that current technologies still cannot generate full 4K accurate visual effects.
@@ 3.4 Information Extraction and Understanding
AI plays a crucial role in automating and optimizing the process of information extraction and understanding, enabling organizations to derive actionable insights from large and diverse data. Yan et al. [87] have categorized information extraction tasks based on the Format-Time-Reference space, as illustrated in Fig. 6 (a), where object detection and video object segmentation (VOS) are considered to be the simplest and the most complex tasks, respectively. Recent advancements in this field draw significant inspiration from LLMs. These advancements include the utilization of prompts as conditional inputs for acquiring information. Moreover, following the pipeline approach used in LLMs, there is a growing trend towards leveraging very large datasets to pre-train models before fine-tuning them for specific downstream tasks. For instance, Meta AI [154] has introduced DINOv2, aimed at enriching information about visual content through self-supervised learning. This model was trained with 142 million carefully selected images, employing the ViT architecture. Google introduced VideoPrism [164], a solution for scene understanding tasks, including classification, localization, retrieval, captioning, and question answering (QA). The model was trained on an extensive and diverse dataset consisting of 36 million high-quality video-text pairs and 582 million video clips accompanied by noisy or machine-generated parallel text.

### v3

@@ 3.3.1 Enhancement
In low-light conditions, scenes often exhibit low contrast, leading to focusing difficulties or the need for long exposures, which can result in blurred images and videos. To address this, LEDNet [279] has introduced a synthetic dataset for such scenarios and incorporated a learnable non-linear activation function within the network to enhance feature intensities. Meanwhile, SNR-Aware [98] estimates spatial-varying Signal-to-Noise Ratio (SNR) maps and proposes local and global learning branches using ResNet and transformer architectures, respectively. NeRCo [108] address low-light problem with INR, which unifies the diverse degradation factors of real-world scenes with a controllable fitting function. Diffusion models (DMs) have also become popular choices for low-light image enhancement [104, 105, 106]. Diff-Retinex [105] formulates the low-light image enhancement problem into Retinex decomposition, and employs multi-path generative diffusion networks to reconstruct the normal-light Retinex probability distribution. A recent state-of-the-art approach presented in [106] decomposes images into high and low frequencies using wavelet transform. High frequencies are enhanced using a transformer-based pipeline, while the low frequencies undergo a diffusion process. This method achieves nearly 2.8dB improvement over the state-of-the-art transformer-based approach, e.g. LLFormer [100], and significantly better than INR-based method, NeRCo [108], on a real low-light image benchmarking dataset. The technique has been extended for video enhancement in [107]. The output of the enhancement typically depends on user preferences. This has been viewed as a one-to-many inverse problem, with attempts to solve it using Bayesian approaches. For example, a Bayesian Enhancement Model (BEM) [280] incorporates Bayesian Neural Networks (BNNs) to capture data uncertainty and produce diverse outputs. The method can be used with Transformers or Mamba as the architecture backbone.
@@
Regarding video enhancement, transformer and DMs are still in their early stages. STA-SUNet [101] has demonstrated that using transformers for low-light video enhancement outperforms CNN-based methods [281]. The recent Mamba-based network [282] also demonstrates promising results, outperforming STA-SUNet by more than 2 dB in PSNR. It is important to note that low-light enhancement is subjective. While most training datasets use normal lighting conditions as ground truth [283], the enhanced images and videos may alter the mood and tone of the content. Therefore, the tools for creative industries should be adjustable, not only for entire images and videos but also adaptive to specific areas and content. For instance, CLE Diffusion [129] enables user-friendly editing of lighting with fine-grained regional controllability.
Recent efforts have focused on enhancing User-Generated Content (UGC) videos—authentic recordings created by individuals rather than brands, often showcasing real experiences with products or services. The winning solution of the NTIRE 2025 Challenge on UGC Video Enhancement [284] implemented a pipeline of four sequential modules: color enhancement, denoising, BasicVSR++ restoration [285], and SwinIR [115]. This method achieved a 17% higher subjective score than the second-place entry, which used a two-stage framework, highlighting a notable improvement in perceived visual quality.
@@ 3.3.2 Style transfer
StyTr2 [109] is the first transformer-based method for style transfer, applying content as a query and style as a key of attention. InST [112] utilizes Stable Diffusion Models as the generative backbone and introduces an attention-based textual inversion module to learn the description of the content. StableVideo [113] uses a text prompt to describe the desired appearance of the output, transforming the input video to have a new look based on a diffusion model. For instance, a video of a white car driving in summer can be altered to show a red car driving in winter. A large pre-trained DM is employed in [111], where the style is injected to manipulate the self-attention of the decoder. To deal with the disharmonious color, they propose an adaptive instance normalization. A survey of style transfer using transformers and diffusion models can be found in [286]. Implicit Neural Representations (INRs) are less commonly used in style transfer tasks due to the difficulty of modeling the cross-representation between style and content. Moon et al. [110] combined INRs with vision transformers for generalizable style transfer; however, the results remain limited in quality. In contrast, the method proposed by Kim et al. [114] uses multilayer perceptrons (MLPs) to map image coordinates to the colors of the stylized output, guided by features extracted from both the content and style inputs to allow controllability.
@@ 3.3.3 Upscaling imagery: super-resolution (SR)
The first use of a transformer, called ESRT, was for capturing long-term dependencies, such as repeating patterns in buildings. This was done in the feature domain extracted by a lightweight CNN module [116], outperforming those that use only CNNs. Since then, most SISR methods have been based on transformers. The Hybrid Attention Transformer (HAT) [118] was introduced, which improves the SR quality over ESRT by more than 2dB when upscaling 2$\times$-4$\times$. However, the NTIRE 2023 Real-Time Super-Resolution Challenge [287] showed that the winner, Bicubic++ [288], uses only convolutional layers and achieves the fastest speed at 1.17ms in upscaling 720p to 4K images. This method is significantly faster than any of the participants in the NTIRE 2025 Challenge [289], where Transformer-based architectures continue to dominate as the mainstream approach.
@@
For video SR, numerous methods have emerged as part of a unified enhancement framework, as discussed in the previous section. One of the pioneering works to incorporate transformers specifically for video SR tasks is the Trajectory-aware Transformer for Video Super-Resolution (TTVSR) [117]. Although the results are slightly inferior to those of BasicVSR++ [285], which employs CNN and was introduced around the same time, both methods significantly enhance detail and sharpness compared to previous approaches, albeit not in real time. To address this limitation, the Deformable Attention Pyramid [291] has been introduced, offering slightly lower quality but a speed-up of over 3$\times$. Recently, Adobe announced their VideoGigaGAN [121], which can perform 8$\times$ upsampling. This is achieved by adding flow estimation and temporal self-attention to the GigaGAN upsampler [120], which is primarily used for image SR, and text-to-image synthesis. Cao et al. [126] introduce a zero-shot video super-resolution framework that leverages a pre-trained image diffusion model, and replaces the spatial self-attention layer with a novel short-long-range (SLR) temporal attention layer. Recently, SeedVR integrated text information (captions) into a Diffusion Transformer (DiT) model, achieving state-of-the-art performance in video super-resolution.
@@ 3.3.4 Restoration
The pioneering transformer-based method for image restoration, SwinIR [115], employs several concatenated Swin Transformer blocks [16]. SwinIR surpasses state-of-the-art CNN-based methods proposed up to the year 2021 in super-resolution and denoising tasks. The model is smaller and reconstructs fine details more effectively. Other two popular approaches that emerged in the same timeframe are Uformer [130] and Restormer [131]. Both incorporate Transformer blocks into hierarchical encoder-decoder networks, employing skip connections similar to those in U-Net. Their objective was to restore noisy images, sharpen blurry images, and remove rain. The networks focused on predicting the residual $R$ and obtaining the restored image $\hat{x}$ through $\hat{x}=y+R$. While their performance is very similar, Restormer has half the parameters of Uformer. More recent, GRL by Li et al. [119] exploits a hierarchy of features in a global, regional, and local range using different ways to compute self-attentions as an image often show similarity within itself in different scales and areas. GRL outperforms SwinIR and Restormer. Additionally, Fei et al. introduced the Generative Diffusion Prior [128] for unsupervised learning, aiming to model posterior distributions for image restoration and enhancement. VmambaIR [143] incorporates Mamba blocks into the U-Net architecture, achieving superior performance compared to SwinIR and Restormer in both visual quality and model size.
@@
For audio restoration, most software discussed in Section 3.1.2 offers tools for enhancing audio quality, such as eliminating background noise, echo, microphone rumble, and occasionally room reverberation, which have been well-established even before the advent of deep learning. There have been efforts to utilize AI for learning global contextual information to aid in the removal of unwanted sounds, leading to better final quality [134]. The latest advancements in this domain are primarily focused on addressing issues where significant portions of the audio data are missing. For instance, Moliner et al. [124] tackle problems such as audio bandwidth extension, inpainting, and declipping by treating them as inverse problems using a diffusion model. For a comprehensive survey on the use of diffusion models in restoration tasks, refer to [292].
@@
i) Deblurring: A lightweight deep CNN model was recently proposed in [293], where a new discriminative temporal feature fusion has been introduced to select the most useful spatial and temporal features from adjacent frames. Feature propagation along the video is done in the wavelet domain. The deblurring performance is comparable to RVRT [99], but it is 5 times faster. DaBiT [133] mitigates focal blur content with depth information and applies SR for further enhancing fine details. Note that not only in software, but AI technologies have also been integrated into hardware. This includes autofocus, which is crucial for capturing sharp images of subjects, especially in dynamic environments where manual adjustments are impractical due to rapid movement. AI-driven autofocus methods have emerged, often tailored for specific camera hardware. For instance, Choi et al. proposed an autofocus model optimized for dual-pixel Canon cameras [294]. Additionally, Yang et al. investigated the correlation between language input and blur map estimation, utilizing semantic cues to enhance autofocus performance [132]. Remarkably, their model achieves comparable results to previous state-of-the-art methods while being more lightweight [295]. Autofocus could be used in conjunction with real-time object tracking (see Section 3.4.3) to produce desirable sharpness for moving objects in the video. Recently, Feng et al. [147] proposed a novel residual diffusion deblurring framework that integrates a conditional diffusion model guided by a defocus map and incorporates residual learning into the single-image defocus deblurring process.
@@
ii) Denoising: SUNet [18] applies Swin transformer blocks combined in a UNet-like architecture. Denoising with diffusion models (DMs) [144] has been proposed by diffusing with estimated noise that is closer to real-world noise rather than Gaussian noise, achieving better performance than SwinIR [115] and Uformer [130]. INR with complex Gabor wavelets as activation functions show promising denoising results [42]. The NTIRE 2025 Image Denoising Challenge [296] revealed that the top-performing methods combined transformer-based and convolutional network architectures. Similarly, recent advances in video denoising also adopt a hybrid approach that integrates both architectures [142, 141].
@@
iii) Dehazing: Vision transformers for single image dehazing was proposed in DehazeFormer [135]. Similar to SUNet, it is UNet-like architecture, but introduced Rescale Layer Normalization for better suit on improving contrast. The Fast Fourier Transform (FFT) has been employed in [140] due to the phase spectrum conveying more structural detail than the amplitude spectrum and demonstrating greater robustness to contrast distortion and noise. Then cross-attention between the RGB and YCbCr color spaces is applied. This approach achieves nearly 5 dB higher PSNR than DehazeFormer on a real-world smoke dataset. For video dehazing, Xu et al. [136] introduced a recurrent multi-range scene radiance recovery module with the space-time deformable attention. They also employs physics prior to inform haze attenuation. This method outperforms DehazeFormer by approximately 1dB.
@@
iv) Mitigating atmospheric turbulence: Similar to dehazing, physics-inspired models have been widely developed to remove turbulence distortion [146, 148], while complex-valued CNNs have been proposed to exploit phase information [297]. There was also an attempt to use instance normalization (INR) to address this issue, providing solutions for tile and blur correction [148]. However, diffusion models have shown superior performance on single-image restoration tasks [145], while transformer-based methods remained the state-of-the-art for video restoration [138, 139]. More recently, Mamba-based architectures have demonstrated their effectiveness in both visual quality and model efficiency [298]. A recent review can be found in [299].
@@ 3.3.5 Inpainting
A comprehensive survey of learning-based image and video inpainting, covering approaches such as CNNs, VAEs, GANs, transformers, and diffusion models, can be found in [302]. Additionally, Elharrouss et al. [303] provide an in-depth review of the current challenges and future directions specific to transformer-based inpainting techniques.
@@ 3.3.6 Image Fusion
Image fusion is the process of merging multiple images from either the same source (such as varying focal points or exposures) or different modalities (e.g. visible and infrared cameras) into a single image. This process integrates complementary information from the various images to enhance overall quality, improve interpretation, and increase the usability of the final image.
Transformers and CNNs have been combined to extract global and local information, respectively. Most methods use CNNs for feature extraction, with transformers operating in the latent space [154, 155]. Notable methods include SwinFusion [154], which utilizes a self-attention-based intra-domain fusion unit and a cross-attention-based inter-domain fusion unit to achieve multi-modal and digital photography image fusion. Transformer-based image fusion has also been applied to downstream tasks like segmentation [156], achieving superior results by leveraging the additional information. Self-attention blocks are employed to enhance intra-feature representations, while the cross-attention mechanism integrates inter-feature information to improve the quality of the fused output [157].
DDFM, the first diffusion model-based image fusion method, estimates noise in the reverse process by combining multiple inputs [158]. The expectation-maximization (EM) algorithm is integrated to estimate the noise distribution parameters, resulting in sharper images compared to traditional DDPM. For an in-depth review, the reader is referred to recent work in [304, 305].
@@ 3.3.7 Editing and Visual Special Effects (VFX)
VFX aims to create and/or manipulate imagery outside the context of a live-action shot in filmmaking and video production. When adding objects, scenes, and effects into traditional photographic videos, generative AI has obviously become an important tool, but some manual operations are still required. For example, in After Effects (EA) [^45] https://www.adobe.com/uk/products/aftereffects.html, the user selects the area where the object will be added and uses text prompts to describe such object. Subsequently, with the current EA version, the user will need to apply motion tracking so the generated object is moved accordingly.
AI technologies can upscale, enhance, and restore low-quality or old footage. For example, standard definition videos can be converted to high definition or even 4K quality without traditional manual remastering processes. This is particularly useful for remastering old movies or enhancing visual details in scenes. Generative AI has also simplified and accelerated automated processes, such as rotoscoping [306], an animation technique where animators trace over motion picture footage frame by frame to create realistic action. AI models can accurately detect and segment objects and characters in video frames, significantly speeding up the post-production process. Additionally, AI can assist the rapid creation of 3D models from 2D images generating realistic animations with minimal input data, facilitating complex human motions and synchronized facial expressions to voiceovers. One restriction is that current technologies still cannot yet generate full 4K accurate visual effects.
@@ 3.4 Information Extraction and Understanding
AI plays a crucial role in automating and optimizing the process of information extraction and understanding, enabling organizations to derive actionable insights from large and diverse data. Yan et al. [91] have categorized information extraction tasks based on the Format-Time-Reference space, as illustrated in Fig. 6 (a), where object detection and video object segmentation (VOS) are considered to be the simplest and the most complex tasks, respectively. Recent advancements in this field draw significant inspiration from LLMs. These advancements include the utilization of prompts as conditional inputs for acquiring information. Moreover, following the pipeline approach used in LLMs, there is a growing trend towards leveraging very large datasets to pre-train models before fine-tuning them for specific downstream tasks. For instance, Meta AI [166] has introduced DINOv2, aimed at enriching information about visual content through self-supervised learning. This model was trained with 142 million carefully selected images, employing the ViT architecture. Google have introduced VideoPrism [178], a tool for scene understanding including classification, localization, retrieval, captioning, and question answering (QA). The model was trained on an extensive and diverse dataset consisting of 36 million high-quality video-text pairs and 582 million video clips accompanied by noisy or machine-generated parallel text.